# Uhlmann and scalar Wilczek-Zee phases of degenerate quantum systems

### v1

@@ Abstract
The Wilczek-Zee (WZ) holonomy arises in degenerate states while the Uhlmann holonomy characterizes finite-temperature topology. We investigate possible relationships between the Uhlmann phase and the scalar WZ phase, which reflects the Uhlmann and WZ holonomy respectively, in an exemplary four-level model with two doubly degenerate subspaces. Through exactly solutions, we contrast the behavior of the Uhlmann and WZ connections and their associated phases. In the zero-temperature limit, the Uhlmann phase may or may not agree with the scalar WZ phase of the degenerate ground states due to the underlying topology. Some conditions are derived in an analytic analysis, under which the Uhlmann phase in the zero-temperature limit converges to the scalar WZ phase. We also briefly discuss possible realizations or simulations of the phenomena as well as their implications.
@@ I Introduction
Previous research has revealed interesting examples where the Uhlmann phase reduces, in the zero-temperature limit, to the Berry phase at zero temperature in non-degenerate systems [31, 35, 41], despite their different origins from topologically distinct principal fiber bundles. Those examples motivate a broader examination of how finite-temperature geometric phases relate to possible zero-temperature counterparts, especially in the presence of degeneracy. Here, we will focus on the Uhlmann phase of degenerate states in the zero-temperature limit. To facilitate a fair comparison with the geometric phase from the WZ holonomy of degenerate ground states, we propose to use the scalar Wilczek-Zee phase, which has a similar definition as the Uhlmann phase to reflect the underlying non-Abelian holonomy, instead of the matrix-valued WZ phases in the search of possible relations with the Uhlmann phase in the zero-temperature limit.
@@
To provide a concrete analysis, we study a four-level model with two doubly degenerate subspaces, which readily exhibits a variety of phenomena for contrasting the behavior of the Uhlmann and scalar WZ phases. Moreover, we explicitly derive the Uhlmann connection and the WZ connection that generate the corresponding non-Abelian holonomy. Through two simple examples of the four-level model, we demonstrate some agreements and disagreements by varying the underlying parameter spaces between the Uhlmann phase in the zero-temperature limit and the scalar WZ phase. An attempt to analytically understand the zero-temperature limit of the Uhlmann holonomy and phase in contrast to the WZ holonomy and its scalar phase further elucidates some interesting but challenging topological properties behind degenerate quantum systems at and beyond zero temperature.
@@
The rest of the paper is organized as follows. Section II introduces the four-level model with two doubly degenerate subspaces and its properties, defines the scalar WZ phase, and summarizes the explicit expressions for the Uhlmann and WZ connections along with their associated holonomy and phases. Section III examines two specific examples of the four-level model to demonstrate possible relationships between the Uhlmann phase in the zero-temperature limit and the scalar WZ phase. Section IV provides a theoretical analysis of the Uhlmann phase in the zero-temperature limit and proposes possible conditions for it to reduce to the scalar WZ phase. Section V discusses some implications of the Uhlmann and scalar WZ phases in possible experiments and simulations. Finally, Section VI summarizes our findings and offers concluding remarks. The Appendix shows some details of the Uhlmann connection.
@@ II.1 The Hamiltonian
where each level is twofold degenerate. We introduce the projectors onto the subspaces with eigenvalues ±$\pm R$ as
@@ II.3 Wilczek-Zee holonomy
for $i$, $j$=$a$, $b$ or $c$, $d$. This definition differs from the conventional one by a factor of $\mathrm{i}=\sqrt{-1}$, but it does not change any physical essence. Thus, the total WZ connection matrix is
@@ II.4 Scalar Wilczek-Zee phase
Since an explicit relation between $\mathcal{A}_{\text{U}}$ in the $T\rightarrow 0$ limit and $\mathcal{A}_{\text{WZ}}$ is difficult to obtain at present, we shift our focus to the Uhlmann and WZ phases arising from the corresponding holonomy. However, the conventional WZ phase is matrix-valued [14], as shown by the elements in Eq. (15). To enable a direct comparison with the zero-temperature limit of the Uhlmann phase, we introduce the scalar Wilczek-Zee phase associated with the degenerate ground-state subspace:
$\theta_{\text{WZ}}(C)=\arg\text{Tr}_{-}\left(\frac{1}{2}P_{-}U_{-}(C)\right),$ (21)
@@
To make the comparison between the $T\rightarrow 0$ Uhlmann phase and scalar WZ phase more transparent, we define the analogue of purification of the ground-state projector as $W_{-}=\frac{1}{\sqrt{D}}\sqrt{P_{-}}\mathcal{U}$. In analogy to the Uhlmann phase in Eq. (13), the scalar WZ phase accumulated along a closed path $C(t)$ ($0\leq t\leq\tau$) can then be expressed as
$\theta_{\text{WZ}}(C)=\arg\langle W_{-}(0)|W_{-}(\tau)\rangle.$ (22)
Here the analogies of the purified state $|W_{-}(t)\rangle$ and the inner products are also introduced in a similar fashion. Since the density matrix approaches $\rho\to\frac{1}{D}P_{-}$ as $T\to 0$, the formally similar expressions of the Uhlmann and scalar WZ phases warrant some investigations into their relations. We set out to search for any correspondence between the Uhlmann phase in the zero-temperature limit and the scalar WZ phase in the four-level model described by Eq. (1) because such a correspondence may serve as a bridge to link the topological properties of mixed states with those of degenerate ground states.
@@ III.1 A simple illustration
Before proceeding further, we note that the behavior of the Uhlmann phase often signals finite-temperature topological phase transitions [31, 34, 35]. At infinitely high temperatures, the density matrix becomes proportional to the identity matrix, implying $\rho\propto 1_{4}$ for a four-level system. As a result, any evolution of $\rho(t)$ along a loop in the parameter space effectively collapses to a single point, thereby rendering the Uhlmann phase trivial [34]. Since $\lim_{T\to 0}\theta_{\text{U}}=\pi$, there must exist a critical temperature $T_{c}$ at which the Uhlmann phase undergoes a sudden jump, indicating a topological phase transition. For the simple case studied here, the critical temperature showing a jump of the Uhlmann phase emerges when $\cos(\pi\chi)=0$ or $\chi=1/2$ since $10<\chi<$ 1 at finite temperatures, which yields $T_{c}\approx 0.75R$. To visualize the transition, we plot $\theta_{\text{U}}$ as a function of $T$ on a semi-logarithmic scale in Figure 1. As shown in the plot, the Uhlmann phase indeed jumps by a quantized value of $\pi$. The change in the Uhlmann phase indicates that below $T_{c}$, the final purification $W(\tau)$ is antiparallel to the initial purification $W(0)$ while above $T_{c}$, $W(\tau)$ becomes parallel to $W(0)$.
@@
This corresponds to the four-level model of Eq. (1) by identifying the components of the vector $\mathbf{R}$ as $R_{1}=\sin k_{x}$, $R_{2}=\sin k_{y}$, $R_{3}=\sin k_{z}$, $R_{4}=\sin k_{u}$, and $R_{5}=m+\cos k_{x}+\cos k_{y}+\cos k_{z}+\cos k_{u}$. The parameter space is the first Brillouin zone with $k_{x}$, $k_{y}$, $k_{z}$, and $k_{u}$ spanning the interval $[-\pi,\pi]$, which forms a four-dimensional torus $T^{4}$.
@@
×$\times\left[(m+3)\cos k_{x}+1\right]\mathrm{d}k_{x},$ (38)
@@
Figure 2: (Top panel) Uhlmann phase $\theta_{\text{U}}$ as a function of temperature $T$ on a semi-logarithmic scale for $m=-3$. A $\pi$-jump occurs at $T\approx 0.75R_{0}$, signaling a topological phase transition. (Bottom panel) Uhlmann-phase diagram of the 4D tight-binding model. The numerical values of $T_{c}$ are marked by the blue circles. The background and dome-shape regions correspond to the trivial ($\theta_{\text{U}}=0$) and topological ($\theta_{\text{U}}=\pi$) phases, respectively. The scalar WZ phase at $T=0$, however, vanishes after the system traverses the path $C$ regardless of $m$. Here $R_{0}=R(m=-3)$.
In the top panel of Fig. 2, we plot the Uhlmann phase $\theta_{\text{U}}$ as a function of temperature $T$ on a semi-logarithmic scale for $m=-3$. A clear $\pi$-jump is observed at the critical temperature $T_{c}=0.75R_{0}$, where $R_{0}=R(m=-3)$. In the bottom panel, we present the Uhlmann-phase diagram of the 4D tight-binding model on the $m\text{-}T$ plane. A finite Uhlmann phase appears exclusively in the range $-4<m<-2$ at $T=0$ and forms a dome-shape region at finite temperatures. The rest of the diagram is occupied by the topologically trivial region with zero Uhlmann phase. By fitting the boundary of the dome-shaped topological region with a symmetric power-law envelope function, we obtain an approximate relation between $T_{c}$ and $m$: $T_{c}/R_{0}\approx 0.75\left[1-(m+3)^{2}\right]^{0.45}$.
@@
in this case. Consequently, the scalar WZ phase vanishes after the loop $C$ is traversed in the parameter space, regardless of the value of $m$. This explicit example demonstrates that the $T\rightarrow 0$ limit of $\theta_{\text{U}}$ does not always reduce to $\theta_{\text{WZ}}$, showcasing a fundamental difference from the previous case. As explained below Eq. (22), the scalar WZ phase is a valid geometric phase of the degenerate ground states, not given simply by the $T\rightarrow 0$ limit of the Uhlmann phase.
@@ IV Correspondence between the Uhlmann and scalar WZ phases
The two examples sharply contrast different outcomes regarding the relationships between the $T\rightarrow 0$ Uhlmann phase and the scalar WZ phase. We now turn to a more general analysis of this issue from an analytic perspective. Our objectives are twofold: to determine whether a simple relationship exists between the Uhlmann connection in the zero-temperature limit and the WZ connection and to identify the conditions under which the Uhlmann phase in the zero-temperature limit may reduce to the scalar WZ phase. In the following, we consider generic systems with $N$-dimensional Hilbert space.
@@
×$\times|\psi^{(i)}_{n}\rangle\langle\psi^{(i)}_{n}|\mathrm{d}\psi^{(j)}_{m}\rangle\langle\psi^{(j)}_{m}|,$ (48)
@@
$\mathcal{A}_{\text{U}}\to$ $-\sum_{i\neq j}\sum_{n=1}^{D_{i}}\sum_{m=1}^{D_{j}}|\psi^{(i)}_{n}\rangle\langle\psi^{(i)}_{n}|\mathrm{d}\psi^{(j)}_{m}\rangle\langle\psi^{(j)}_{m}|$
$=$ $-\sum_{i,j}\sum_{n=1}^{D_{i}}\sum_{m=1}^{D_{j}}|\psi^{(i)}_{n}\rangle\langle\psi^{(i)}_{n}|\mathrm{d}\psi^{(j)}_{m}\rangle\langle\psi^{(j)}_{m}|$
$+\sum_{i}\sum_{n,m=1}^{D_{i}}|\psi^{(i)}_{n}\rangle\langle\psi^{(i)}_{n}|\mathrm{d}\psi^{(i)}_{m}\rangle\langle\psi^{(i)}_{m}|$
@@
$\mathcal{A}_{\text{WZ}}=$ $\sum_{i}\sum_{n,m=1}^{D_{i}}\langle\psi^{(i)}_{n}|\mathrm{d}\psi^{(i)}_{m}\rangle|\psi^{(i)}_{n}\rangle\langle\psi^{(i)}_{m}|$
@@
is the full WZ connection matrix. Here we have suppressed the explicit dependence on the parameter $\mathbf{R}$ for brevity. According to Eq. (IV), the Uhlmann holonomy along a closed loop $C$ in the parameter space is expressed as:
@@
In the first example above, when the system evolves along the equator, the eigenvalues ±$\pm R$ remain constant, and the associated evolution corresponds to a unitary transformation of the states involved. This can be written as $|\psi^{(j)}_{m}(\mathbf{R}(t))\rangle=\mathcal{D}(\mathbf{R}(t))|\psi^{(j)}_{m}(\mathbf{R}(0))\rangle$. For notational simplicity, we denote $|\psi^{(j)}_{m}(0)\rangle=|\psi^{(j)}_{m}(\mathbf{R}(0))\rangle$. In this scenario, the relationship between $\mathcal{A}_{\text{U}}$ in the $T\rightarrow 0$ limit and $\mathcal{A}_{\text{WZ}}$ simplifies to
@@
Assuming the additional condition $[\mathcal{A}_{\text{WZ}},\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}]=0$ holds, the expression further reduces to
@@
Since $\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}$ is anti-Hermitian, the path-ordered exponentiation $\mathcal{P}\mathrm{e}^{{\oint}_{C}\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}}$ forms an $N\times N$ unitary matrix. Given that this path-ordered exponentiation maps the closed loop $C\sim S^{1}$ to the group $U(N)$ of unitary matrices, the resulting $\mathcal{P}\mathrm{e}^{{\oint}_{C}\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}}$ is an element of the fundamental group $\pi_{1}(U(N))=\mathbb{Z}$. Thus, we obtain [44]

### v2

@@ Abstract
The Wilczek-Zee (WZ) holonomy arises in degenerate states while the Uhlmann holonomy characterizes finite-temperature topology. We investigate possible relationships between the Uhlmann phase and the scalar WZ phase, which reflects the Uhlmann and WZ holonomy respectively, in an exemplary four-level model with two doubly degenerate subspaces. Through exact solutions, we contrast the behavior of the Uhlmann and WZ connections and their associated phases. In the zero-temperature limit, the Uhlmann phase may or may not agree with the scalar WZ phase of the degenerate ground states due to obstructions from the Hamiltonian manifested as Dirac points. This is in stark contrast to non-degenerate systems where the correspondence between the Uhlmann and Berry phases in general holds. Our analyses further show that for the example studied here, the Uhlmann phase catches the singular behavior at the Dirac points while the WZ connection and scalar WZ phase vanish along a zero-field axis. We also briefly discuss possible experimental implications.
@@ I Introduction
Previous research has revealed interesting examples where the Uhlmann phase reduces, in the zero-temperature limit, to the Berry phase at zero temperature in non-degenerate systems [31, 35, 41], despite their different origins from topologically distinct principal fiber bundles. By presenting a concrete example that exhibits a violation of the correspondence for degenerate cases, we will identify topological obstructions due to the presence of Dirac points when the Hamiltonian vanishes and the system becomes completely degenerate. This dichotomy motivates a broader examination of how finite-temperature geometric phases relate to possible zero-temperature counterparts, especially in the presence of degeneracy. However, it may not be meaningful to directly compare the Uhlmann phase with the matrix-valued WZ phase in the presence of degeneracy. To facilitate a fair comparison with the geometric phase from the WZ holonomy of degenerate ground states, we propose to use the scalar Wilczek-Zee phase, which has a similar definition as the Uhlmann phase to reflect the underlying non-Abelian holonomy, instead of the matrix-valued WZ phases in the search of possible relations with the Uhlmann phase in the zero-temperature limit.
@@
To provide a concrete analysis, we study a four-level model with two doubly degenerate subspaces, which readily exhibits a variety of phenomena for contrasting the behavior of the Uhlmann and scalar WZ phases. Moreover, we explicitly derive the Uhlmann connection and the WZ connection that generate the corresponding non-Abelian holonomy. Through two simple examples of the four-level model, we demonstrate some agreements and disagreements by varying the underlying parameter spaces between the Uhlmann phase in the zero-temperature limit and the scalar WZ phase. An attempt to analytically understand the zero-temperature limit of the Uhlmann holonomy and phase in contrast to the WZ holonomy and its scalar phase further elucidates some interesting but challenging topological properties behind degenerate quantum systems at and beyond zero temperature, particularly in the presence of singular behavior from the Hamiltonian.
@@
The rest of the paper is organized as follows. Section II introduces the four-level model with two doubly degenerate subspaces and its properties, defines the scalar WZ phase, and summarizes the explicit expressions for the Uhlmann and WZ connections along with their associated holonomy and phases. Section III examines two specific examples of the four-level model to demonstrate possible relationships between the Uhlmann phase in the zero-temperature limit and the scalar WZ phase. Section IV provides a theoretical analysis of the Uhlmann phase in the zero-temperature limit and proposes possible conditions for reduction to the scalar WZ phase in degenerate systems. Section V analyzes the existence of a zero-field axis where the WZ connection vanishes and the Uhlmann connection near a Dirac point in the 4D tight-binding model to pinpoint the difference between the zero-temperature limit of the Uhlmann phase and the scalar WZ phase. Section VI discusses some implications in possible experiments and simulations. Finally, Section VII concludes our work. The Appendix summarizes some details and subtleties in the definitions and evaluations of various geometric phases, presents a geometric proof of the zero-field axis condition, along with a conditional proof of the Uhlmann-Berry correspondence for non-degenerate systems.
@@ II.1 The Hamiltonian
where each level is twofold degenerate. We introduce the projectors onto the subspaces with eigenvalues $\pm R$ as
@@ II.3 Wilczek-Zee holonomy
for $i$, jj=$a$, $b$ or $c$, $d$. This definition differs from the conventional one by a factor of $\mathrm{i}=\sqrt{-1}$, but it does not change any physical essence. Thus, the total WZ connection matrix is
@@ II.4 Scalar Wilczek-Zee phase
Since an explicit relation between $\mathcal{A}_{\text{U}}$ in the $T\rightarrow 0$ limit and $\mathcal{A}_{\text{WZ}}$ is difficult to obtain in general, we shift our focus to the Uhlmann and WZ phases arising from the corresponding holonomy. However, the conventional WZ phase is matrix-valued [14], as shown by the elements in Eq. (15). To enable a direct comparison with the zero-temperature limit of the Uhlmann phase, we introduce the scalar Wilczek-Zee phase associated with the degenerate ground-state subspace:
$\theta_{\text{WZ}}(C)=\arg\text{Tr}_{-}\left(\frac{1}{2}P_{-}(0)U_{-}(C)\right),$ (21)
@@
We remark that the Uhlmann and scalar WZ phases represent the phase of the average of the holonomy over the density matrix and degenerate states, respectively. The two scalar phases facilitate a direct comparison despite the different underlying holonomy structures. We caution that when a zero scalar phase arises from a nontrivial holonomy, it implies the averaged holonomy leads to a positive real number and thus does not contribute to the phase.
To make the comparison between the $T\rightarrow 0$ Uhlmann phase and scalar WZ phase more transparent, we define the analogue of purification of the ground-state projector as $W_{-}=\frac{1}{\sqrt{D}}\sqrt{P_{-}}\mathcal{U}$, where $\mathcal{U}$ represents the phase factor associated with the degenerate ground-state subspace of dimension $D$. Under cyclic evolution along a closed path $C(t)$ ($0\leq t\leq\tau$) in the parameter space, the initial and final phase factors differ by a WZ holonomy:
$\mathcal{U}(\tau)=\mathcal{P}\mathrm{e}^{-{\oint}_{C}\mathcal{A}_{-}}\mathcal{U}(0)=U_{-}(C)\mathcal{U}(0),$ (22)
which is the counterpart of the Uhlmann holonomy in Eq. (8). In a direct analogy to the Uhlmann phase defined in Eq. (13), we express the scalar WZ phase as:
$\theta_{\text{WZ}}(C)=\arg\langle W_{-}(0)|W_{-}(\tau)\rangle,$ (23)
which exactly coincides with Eq. (21) (see Appendix B for the consistency of the definitions). Here the analogies of the purified state $|W_{-}(t)\rangle$ and the inner products are also introduced in a similar fashion. Since the density matrix approaches $\rho\to\frac{1}{D}P_{-}$ as $T\to 0$, the formally similar expressions of the Uhlmann and scalar WZ phases warrant some investigations into their relations. We set out to search for any correspondence between the Uhlmann phase in the zero-temperature limit and the scalar WZ phase in the four-level model described by Eq. (1) because such a correspondence may serve as a bridge to link the topological properties of mixed states with those of degenerate ground states.
@@ III.1 A simple illustration
Before proceeding further, we note that the Uhlmann phase can signal finite-temperature topological phase transitions [31, 34, 35] with quantized jumps. At infinitely high temperatures, the density matrix becomes proportional to the identity matrix, implying $\rho\propto 1_{4}$ for a four-level system. As a result, any evolution of $\rho(t)$ along a loop in the parameter space effectively collapses to a single point, thereby rendering the Uhlmann phase trivial [34]. Since $\lim_{T\to 0}\theta_{\text{U}}=\pi$, there must exist a critical temperature $T_{c}$ at which the Uhlmann phase undergoes a sudden jump, indicating a temperature-induced topological phase transition. For the simple case studied here, the critical temperature showing a jump of the Uhlmann phase emerges when $\cos(\pi\chi)=0$ or $\chi=1/2$ since $0<\chi<1$ at finite temperatures, which yields $T_{c}\approx 0.75R$.
To visualize the transition, we plot $\theta_{\text{U}}$ as a function of $T$ on a semi-logarithmic scale in Figure 1. As shown in the plot, the Uhlmann phase indeed jumps by a quantized value of $\pi$. The quantized jump in the Uhlmann phase, which signifies a temperature-induced topological phase transition, indicates that below $T_{c}$, the final purification $W(\tau)$ is antiparallel to the initial purification $W(0)$ while above $T_{c}$, $W(\tau)$ becomes parallel to $W(0)$. Such a change of the topological property resembles that between a Mobius strip and a cylinder [43]. Therefore, the temperature-induced topological phase transition shown in Figure 1 has the Hamiltonian and eigenstates intact as temperature varies, so it is not a typical phase or topological transition due to a change of thermodynamics or ground states but features a change of the topological object (the holonomy) in the underlying geometry of the mixed states. We notice the number of critical temperatures where the Uhlmann phase exhibits quantized jumps may increase with the dimension of the Hilbert space, as illustrated in Ref. [34]. However, the number of critical temperatures is not expected to increase in the presence of degeneracy.
@@
This corresponds to the four-level model of Eq. (1) by identifying the components of the vector $\mathbf{R}$ as $R_{1}=\sin k_{x}$, $R_{2}=\sin k_{y}$, $R_{3}=\sin k_{z}$, $R_{4}=\sin k_{u}$, and $R_{5}=m+\cos k_{x}+\cos k_{y}+\cos k_{z}+\cos k_{u}$. The parameter space is the first Brillouin zone with $k_{x}$, $k_{y}$, $k_{z}$, and $k_{u}$ spanning the interval $\pi,\pi]$, which forms a four-dimensional torus $T^{4}$.
@@
$\times\left[(m+3)\cos k_{x}+1\right]\mathrm{d}k_{x},$ (39)
@@
Figure 2: (Top panel) Uhlmann phase $\theta_{\text{U}}$ as a function of temperature $T$ on a semi-logarithmic scale for m=3m=-3. A $\pi$-jump occurs at $T\approx 0.75R_{0}$, signaling a temperature-induced topological phase transition. (Bottom panel) Uhlmann-phase diagram of the 4D tight-binding model. The numerical values of $T_{c}$ are marked by the blue circles. The background and dome-shape regions correspond to the trivial ($\theta_{\text{U}}=0$) and topological ($\theta_{\text{U}}=\pi$) phases, respectively. The scalar WZ phase at $T=0$, however, vanishes after the system traverses the path $C$ regardless of $m$. Here $R_{0}=R(m=-3)$.
In the top panel of Fig. 2, we plot the Uhlmann phase $\theta_{\text{U}}$ as a function of temperature $T$ on a semi-logarithmic scale for m=3m=-3. A clear $\pi$-jump is observed at the critical temperature $T_{c}=0.75R_{0}$, where $R_{0}=R(m=-3)$. In the bottom panel, we present the Uhlmann-phase diagram of the 4D tight-binding model on the $m\text{-}T$ plane. A finite Uhlmann phase appears exclusively in the range 4<m<2-4<m<-2 at $T=0$ and forms a dome-shape region at finite temperatures. The rest of the diagram is occupied by the topologically trivial region with zero Uhlmann phase. By fitting the boundary of the dome-shaped topological region with a symmetric power-law envelope function, we obtain an approximate relation between $T_{c}$ and $m$: $T_{c}/R_{0}\approx 0.75\left[1-(m+3)^{2}\right]^{0.45}$.
@@
in this case. Consequently, the scalar WZ phase vanishes after the loop $C$ is traversed in the parameter space, regardless of the value of $m$. This explicit example demonstrates that the $T\rightarrow 0$ limit of $\theta_{\text{U}}$ does not always reduce to $\theta_{\text{WZ}}$, showcasing a fundamental difference from the previous case. As explained below Eq. (23), the scalar WZ phase is a valid geometric phase of the degenerate ground states, not given simply by the $T\rightarrow 0$ limit of the Uhlmann phase. We will analyze and explain the disagreement in the 4D tight-binding model in later discussions.
@@ IV Correspondence between the Uhlmann and scalar WZ phases
Here we first review the correspondence between the $T\rightarrow 0$ Uhlmann phase and Berry phase for non-degenerate systems. In an attempt to develop a similar correspondence, we found obstructions from degeneracy, symmetry, and topological properties exemplified by the 4D tight-binding model.
IV.1 Complications from degeneracy
For generic non-degenerate cases, a conditional proof of the correspondence between the $T\rightarrow 0$ Uhlmann phase and the ground-state Berry phase has been given in Ref. [41]. Appendix C presents a more general proof with less stringent conditions. The validity of the correspondence hinges on two conditions (adiabaticity and non-degeneracy) that are violated in systems exhibiting Dirac points, which represent critical points in the parameter space where the energy gap of the system collapses (for example, $R\to 0$ in Eq. (1)), resulting in degeneracy. These singularities emerge at energy-level crossings and may lead to non-analytic behavior in geometric observables.
These considerations render the correspondence between the Uhlmann phase in the zero-temperature limit and the Berry phase inapplicable to systems with degeneracy and Dirac points. We will present a more careful analysis and examine how topological obstructions, including Dirac points due to the Hamiltonian and zero-field axis due to symmetry in the 4D tight-binding model, prevent an agreement between the Uhlmann phase in the zero-temperature limit and the geometric phase of the ground state when degeneracy is present.
IV.2 Correspondence between Uhlmann and WZ phases is not guaranteed
The two examples discussed in Sec. III sharply contrast different outcomes regarding the relationships between the $T\rightarrow 0$ Uhlmann phase and the scalar WZ phase. We now turn to a more general analysis of this issue from an analytic perspective. Our objectives are twofold: to determine whether a simple relationship exists between the Uhlmann connection in the zero-temperature limit and the WZ connection and to identify the conditions under which the Uhlmann phase in the zero-temperature limit may reduce to the scalar WZ phase. In the following, we consider generic systems with $N$-dimensional Hilbert space.
@@ IV.2 Correspondence between Uhlmann and WZ phases is not guaranteed
[…]
@@
[…]
@@
[…]
@@
is the full WZ connection matrix. Here we have suppressed the explicit dependence on the parameter $\mathbf{R}$ for brevity. According to Eq. (IV.2), the Uhlmann holonomy along a closed loop $C$ in the parameter space is expressed as:
@@
In the first example above, when the system evolves along the equator, the eigenvalues $\pm R$ remain constant, and the associated evolution corresponds to a unitary transformation of the states involved. This can be written as $|\psi^{(j)}_{m}(\mathbf{R}(t))\rangle=\mathcal{D}(\mathbf{R}(t))|\psi^{(j)}_{m}(\mathbf{R}(0))\rangle$. For notational simplicity, we denote $|\psi^{(j)}_{m}(0)\rangle=|\psi^{(j)}_{m}(\mathbf{R}(0))\rangle$. In this scenario, the relationship between $\mathcal{A}_{\text{U}}$ in the $T\rightarrow 0$ limit and $\mathcal{A}_{\text{WZ}}$ simplifies to
@@
Assuming the additional condition $\mathcal{A}_{\text{WZ}},\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}]=0$ holds, the expression further reduces to
@@
Since $\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}$ is anti-Hermitian, the path-ordered exponentiation $\mathcal{P}\mathrm{e}^{{\oint}_{C}\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}}$ forms an $N\times N$ unitary matrix. Given that this path-ordered exponentiation maps the closed loop $C\sim S^{1}$ to the group $U(N)$ of unitary matrices, the resulting $\mathcal{P}\mathrm{e}^{{\oint}_{C}\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}}$ is an element of the fundamental group $\pi_{1}(U(N))=\mathbb{Z}$. Moreover, since $\mathrm{d}\mathcal{D}\mathcal{D}^{\dagger}$ is the Maurer-Cartan form of the $U(N)$ group, we then obtain [45]